Write the equation of the line in slope-intercept form that passes through the point (3, -2) and has a slope of 2.
y = 2x - 8
y = 2x - 4
y = 2x + 7
y = 2x + 5

Answer by mananth(16946)

y-y1=m(x-x1)
y-(-2)=2(x-3)
y+2=2x-6
y = 2x - 8